A monopolist produces a good that consumers demand according to the function P (Q) = 226 - 1/2Q. The firm's total cost function is TC (Q) = 42Q + 12, 500, with constant marginal cost MC (Q) = 42. Initially, the firm maximizes their profits by charging one price to consumers. The firm will produce output, where each good will be sold for $ units of Later, the firm practices perfect first-degree price discrimination. The firm's new profit-maximizing output level is The firm will earn total revenue equal to $ Compared to the single-price monopolist equilibrium, the amount of DWL in the market will when the monopolist practices perfect first-degree price-discrimination. Choose from: increase, decrease, remain the same. Make sure you are typing in your answer exactly as provided in the answer options. Next


1
Expert's answer
2021-12-22T16:03:08-0500

The rule of profit maximization by a monopolist

MR=MC<P

MC=42

P (Q) = 226 - 1/2Q

MR = TR’(Q)=(226-0.5Q)Q=226-Q

226-Q=42

Q=184

"P=226-0.5\times184=226-92=134"

Condition for maximizing monopoly revenue: MR = 0

226-Q=0

Q=226

"πmax = TR – TC=134\times184-(42\times184+12500)=24656-7728-12500=4428"

Perfect price discrimination is the first degree when a monopolist in the market receives all the surplus of the consumer. The product is sold to all segments of Central Asia at the cost that they are "more willing to pay than refuse." Also, the policy of price discrimination for this type is called differentiation by client income.

The first price discrimination cannot be applied in real conditions. The fact is that the appointment of a separate price for each buyer is possible only under ideal conditions. In addition, the company, as a rule, is not able to determine the exact cost that each consumer is willing to pay.

A monopoly generates fewer surpluses and is less efficient than a competitive market, and therefore leads to irretrievable losses. Monopoly pricing creates irretrievable losses because the firm refuses to deal with consumers.

The reorganization of a completely competitive industry into a monopoly leads to irretrievable losses for society. He also transfers to the monopoly firm a part of the consumer's surplus earned in the event of competition.


increase